OpenStudy (anonymous):

Given tan theta=-5/12 and theta is in Quadrant IV. cot theta=? a. -13/5 b. -12/5 c. 12/5 d. 13/5

OpenStudy (anonymous):

flip it

OpenStudy (anonymous):

\[\cot(\theta)=\frac{1}{\tan(\theta)}\]

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

yw